TK. 10,000 is invest for 4 years at 5% interest. What will be the difference between simple interest and compound interest in the case?


please answer



1
Expert's answer
2022-02-14T17:24:33-0500

Amount, "A=P(1+r/100)^n=10000(1+5/100)^4=\$12155.0625"

Interest "=12155.0625-10000=\$2155.0625"


Simple interest, "I=\frac{PRT}{100}=\frac{10000\times 5 \times 4}{100}=\$2000"

Difference ="2155.0625-2000=\$155.0625"